### v2

@@ Abstract
Video-to-video editing involves editing a source video along with additional control (such as text prompts, subjects, or styles) to generate a new video that aligns with the source video and the provided control. Traditional methods have been constrained to certain editing types, limiting their ability to meet the wide range of user demands. In this paper, we introduce AnyV2V, a novel training-free framework designed to simplify video editing into two primary steps: (1) employing an off-the-shelf image editing model (e.g. InstructPix2Pix, InstantID, etc) to modify the first frame, (2) utilizing an existing image-to-video generation model (e.g. I2VGen-XL) for DDIM inversion and feature injection. In the first stage, AnyV2V can plug in any existing image editing tools to support an extensive array of video editing tasks. Beyond the traditional prompt-based editing methods [20, 55, 15, 23, 36, 54], AnyV2V also can support novel video editing tasks, including reference-based style transfer, subject-driven editing, and identity manipulation, which were unattainable by previous methods. In the second stage, AnyV2V can plug in any existing image-to-video models to perform DDIM inversion and intermediate feature injection to maintain the appearance and motion consistency with the source video. On the prompt-based editing, we show that AnyV2V can outperform the previous best approach by 35% on prompt alignment, and 25% on human preference. On the three novel tasks, we show that AnyV2V also achieves a high success rate. We believe AnyV2V will continue to thrive due to its ability to seamlessly integrate the fast-evolving image editing methods. Such compatibility can help AnyV2V to increase its versatility to cater to diverse user demands.
Keywords:
video editing diffusion models plug-and-play
Figure 1: We build a universal framework AnyV2V to handle all types of video-to-video editing tasks conditioned on different inputs. AnyV2V disentangles the video editing process into two stages: (1) first-frame image editing and (2) image-to-video generation with DDIM inversion + feature injection.
@@ 1 Introduction
The task of video-to-video editing represents a crucial area of research that enables users to edit and create videos according to their preferences seamlessly. In this domain, an AI model is tasked with processing a source video along with various forms of guidance—such as text prompts, facial identity, styles, subjects, depth, and more—to synthesize a new video. It is imperative that the synthesized video not only remains faithful to the source video but also accurately incorporates the additional guidance provided. This requirement ensures that the synthesized video meets the specific desires and intentions of the user. In order to tackle the video-to-video generation tasks, an extensive range of methods [20, 15, 54, 55, 3, 36, 35, 23, 15] have been proposed. These methods are built on top of the open-source image or video diffusion models [45, 17, 62, 9, 53]. However, these methods are all limited to performing specific types of editing tasks. For instance, TokenFlow [20] excels at text-guided style transfer but falls short in localized editing. VideoP2P [36] are restricted to follow certain types of editing prompts. Besides, some of these approaches [55, 23, 17] require additional parameter tuning or video feature extraction (e.g. tracking, point correspondence, etc) to ensure appearance or temporal consistency. Such overhead could lead to much higher time and memory costs. These limitations make them incapable of meeting the diverse user requirements, necessitating the establishment of separate video editing workflows for different tasks.
In recognition of these limitations, a compelling need arises to conceptualize and develop a simple and general-purpose video editing solution. In this paper, we introduce a unified yet simple framework, AnyV2V, designed to address (possibly) any video editing tasks. Our primary insight is that any video editing task can be effectively decomposed into two pivotal stages:
@@
    Apply an off-the-shelf task-specific image editing model over the first frame.
@@
    Apply an image-to-video model to perform DDIM inversion on the source video and then inject the intermediate features to generate the new video.
@@
In the first stage, AnyV2V employs an existing task-specific image editing method to edit the first frame(Section 4.1). In the second stage, we aim to propagate the edited effect across the entire video while ensuring alignment with the source video. To achieve this, we follow an invert-and-generate framework [56, 38]. In contrast to prior methods that employed text-to-video models [3, 4] for DDIM inversion [47], we employ the image-to-video models [62, 12, 44] for DDIM inversion(Section 4.2) to enable the first-frame condition. With the inverted latents as initial noise and the modified first frame as the conditional signal, the I2V model is able to generate videos that are not only faithful to the edited first frame but also follow the appearance and motion of the source video. To further enforce the consistency of the appearance and motion with the source video, we perform feature injection in convolution layers, spatial attention layers and temporal attention layers in the denoising U-Net of the I2V model. These techniques are discussed in detail in Section 4.3 and Section 4.4, and we demonstrate their effectiveness by conducting an ablation study in Section 5.5. By performing the two-stage editing process, AnyV2V effectively offloads the editing operation to existing image editing tools. This disentanglement helps AnyV2V excel in:
  1. 1.
    Compatibility: AnyV2V has superior compatibility with all the image editing methods. We demonstrate that AnyV2V can seamlessly build on top of advanced image editing methods, such as InstructPix2Pix [6], InstantID [51], NST [18], AnyDoor [11] to perform diverse types of editing.
  2. 2.
@@
    Simplicity: AnyV2V is a tuning-free approach without requiring any additional video features to achieve high appearance and temporal consistency.
In this paper, we showcase AnyV2V’s versatility and effectiveness by the qualitative and quantitative results. We comprehensively evaluate our method on four important tasks: (1) prompt-based editing, (2) reference-based style transfer, (3) subject-driven editing and (4) identity manipulation. We collected a new dataset of roughly 100 examples based on the four tasks to evaluate the performance of AnyV2V and other baselines [20, 15, 55]. We show that AnyV2V is among the first to perform (2) reference-based style transfer, (3) subject-driven editing, and (4) identity manipulation in the video domain. Moreover, the AnyV2V exhibits 25% human preference improvement over the previous best model [20] in (1) prompt-based editing. We also perform a comprehensive ablation study to show the impact of our design choices. Our contributions are summarized as follows:
  1. 1.
    We introduce AnyV2V, a plug-and-play unified framework tailored for a diverse range of video-to-video editing tasks.
  2. 2.
    To the best of our knowledge, we are the first to perform video editing using pre-trained I2V models, marking a novel paradigm in this field.
  3. 3.
    Universal compatibility with image editing methods. AnyV2V is compatible with any image editing method, extending any image editing method into the video domain at no cost.
  4. 4.
    We show, both quantitatively and qualitatively, that our method outperforms existing state-of-the-art baselines on prompt-based editing and exhibits robust performance on three novel tasks: reference-based style transfer, subject-driven editing, and identity manipulation.
2 Related Work
2.0.1 Text-to-Video (T2V) Generation.
Recent advancements in diffusion models [27, 46] have led to a huge development in text-to-video (T2V) generation. Existing diffusion-based T2V models generally follow the paradigm of extending text-to-image (T2I) generation models to the video domain, where additional temporal modelling layers are added to the T2I model to enable video generation. Recent T2V generation models are developed from pixel-space diffusion models [29, 26, 19], latent diffusion models (LDMs) [63, 1, 5, 24, 8, 53], as well as diffusion transformers [25, 10, 41]. LDM-based [45] T2V generation models are the most extensively studied methods due to their high efficiency and the open-sourced Stable Diffusion [45] models. T2V models have been shown to be foundational models for a wide range of applications such as I2V generation [57], personalized video generation [39] and video editing [3].
2.0.2 Image-to-Video (I2V) Generation.
To achieve more precise control over the video generation process, a series of I2V generation methods employing additional image conditions in the video generation process have emerged. Earlier efforts such as VideoComposer [52] and VideoCrafter1 [8] use the input image as a style reference and cannot fully preserve the visual details in the output videos. More recently, methods such as DynamiCrafter [57], I2VGen-XL [62], SEINE [12] and ConsistI2V [44] have overcome this problem and can generate videos that are more consistent with the input image (often used as the first frame of the video). I2V generation models have also been employed in a wide range of applications such as long video generation [12, 59, 44], generative frame interpolation [57, 12] and video storytelling [57, 59]. However, to the best of our knowledge, no methods have applied I2V generation models for video editing. We aim to investigate this area and propose a highly flexible training-free framework for video editing using I2V generation models.
2.0.3 Semantic Image and Video Editing.
Visual generation has attracted considerable attention within the field; however, visual manipulation also represents a significant and popular area of interest. The research area of image editing encompasses a wide range of tasks. The popular tasks include stylization [18, 13, 37], identity manipulation [51], subject-driven editing [34, 22, 11], and localized editing [40, 2, 48, 33]. One interesting direction is to train a model to understand text inputs to perform general image editing tasks [60, 6, 49]. However, image editing remained an open and challenging problem as none of the work was able to tackle all the tasks effectively [31]. On the other hand, the video editing problem was often perceived as an extension of image editing problems, incorporating the time dimension. Recent video editing works [20, 54, 42, 15, 58, 7] employ text prompts to execute video editing tasks. These approaches have opened new avenues for interactive and intuitive video manipulation, allowing users to guide the editing process through natural language. However, such methodologies fail to offer precise control to the user, often resulting in edits that may not fully align with the user’s intentions or the desired level of detail due to the ambiguity of natural language and the limit of model capability. Our work is the first to enable precise video editing using pre-trained Image-to-Video (I2V) models. More importantly, our paradigm merges the two problems into a single domain to create a unified solution, offering high controllability in video editing tasks.
3 Preliminary
3.0.1 Image-to-Video (I2V) Generation Models.
$Q=W^{Q}z,K=W^{K}z,V=W^{V}z,$ (1)
3.0.2 DDIM Inversion.
3.0.3 Plug-and-Play (PnP) Diffusion Features.
@@ 4 AnyV2V
Figure 2: AnyV2V framework. Our framework takes a source video $V^{S}$ as input. In the first stage, we apply a block-box image editing method on the first frame $I_{1}$ according to the editing task. In the second stage, the source video is inverted to initial noise $z_{T}^{S}$, which is then denoised using DDIM sampling. During the sampling process, we extract spatial features, spatial attention, and temporal attention from the image-to-video models’ decoder layers. To generate our edited video, we perform a DDIM sampling by fixing $z_{T}^{*}$ as $z_{T}^{T}$ and use the edited first frame as the conditional signal. During the sampling, we inject the features and attention into corresponding layers of the model.
Our method presents a two-stage approach to video editing. Given a source video $V^{S}=\{I_{1},I_{2},I_{3},...,I_{n}\}$, where $I_{i}$ is the frame at time $i$ and $n$ dentes the video length. We extract the initial frame $I_{1}$ and pass it into an image editing model $\phi_{\text{img}}$ to obtain an edited first frame $I^{*}_{1}=\phi_{\text{img}}(I_{1},C)$, where $C$ is the auxiliary conditions for image editing models such as text prompt, mask, style, etc. In the second stage, we feed the edited first frame $I^{*}_{1}$ and a target prompt $\mathbf{s}^{*}$ into an I2V generation model $\epsilon_{\theta}$ and employ the inverted latent from the source video $V^{S}$ to guide the generation process such that the edited video $V^{*}$ follows the motion of the source video $V^{S}$ and the semantic information represented in the edited first frame $I^{*}_{1}$ and the target prompt $\mathbf{s^{*}}$. An overall illustration of our video editing pipeline is shown in Figure 2. In this section, we explain each core component of our method.
@@ 4.1 Flexible First Frame Editing
In visual manipulation, controllability is a key element in performing precise editing. Current video editing models lack this controllability as most of the existing methods can only edit videos based on target text prompts (e.g. TokenFlow [20]) or text instructions (e.g. InsV2V [14]). Our AnyV2V enables more controllable video editing by utilizing image editing models to modify the video’s first frame. This strategic approach enables highly accurate modifications in the video and is compatible with a broad spectrum of image editing models, including other deep learning models that can perform image style transfer [18, 21, 37], mask-based image editing [40, 2], image inpainting [48, 33], identity-preserving image editing [51], and subject-driven image editing [11, 34, 22]. This stage is highly flexible and it can even be done by human experts.
@@ 4.3 Appearance Guidance via Spatial Feature Injection
Our empirical observation (cf. Section 5.5) suggests that I2V generation models already have some editing capabilities by only using the edited first frame and DDIM inverted noise as the model input. However, we find that this simple approach is often unable to correctly preserve the background in the edited first frame and the motion in the source video, as the conditional signal from the source video encoded in the inverted noise is limited.
To enforce consistency with the source video, we perform feature injection in both convolution layers and spatial attention layers in the denoising U-Net. During the video sampling process, we simultaneously denoise the source video using the previously collected DDIM inverted latents $\mathbf{z}^{S}_{t}$ at each time step $t$ such that $\mathbf{z}^{S}_{t-1}=\epsilon_{\theta}(\mathbf{z}^{S}_{t},I_{1},\varnothing,t)$. We preserve two types of hidden features during source video denoising: convolution features $f^{l_{1}}$ before skip connection from the $l_{1}^{\text{th}}$ residual block in the U-Net decoder, and the spatial self-attention scores $\{A_{s}^{l_{2}}\}$ from $l_{2}=\{l_{low},l_{low+1},...,l_{high}\}$ layers. We collect the queries $\{Q_{s}^{l_{2}}\}$ and keys $\{K_{s}^{l_{2}}\}$ instead of directly collecting $A_{s}^{l_{2}}$ as the attention score matrices are parameterized by the query and key vectors. We then replace the corresponding features during denoising the edited video in both the normal denoising branch and the negative prompt branch for classifier-free guidance [28]. We use two thresholds $\tau_{conv}$ and $\tau_{sa}$ to control the convolution and spatial attention injection to only happen in the first $\tau_{conv}$ and $\tau_{sa}$ steps during video sampling.
@@ 4.4 Motion Guidance through Temporal Feature Injection
To better reconstruct the source video motion in the edited video, we propose to also inject the temporal attention features in the video generation process. Similar to spatial attention injection, we collect the source video temporal self-attention queries $Q^{l_{3}}_{t}$ and keys $K^{l_{3}}_{t}$ from some U-Net decoder layers represented by $l_{3}$ and inject them into the edited video denoising branches. We also only apply temporal attention injection in the first $\tau_{ta}$ steps during sampling. Overall, combining the spatial and temporal feature injection mechanisms, the denoising process of our AnyV2V can be represented by:
$\mathbf{z}^{*}_{t-1}=\epsilon_{\theta}(\mathbf{z}^{*}_{t},I^{*},\mathbf{s}^{*},t,\{f^{l_{1}},Q_{s}^{l_{2}},K_{s}^{l_{2}},Q_{t}^{l_{3}},K_{t}^{l_{3}}\}).$ (4)
Our proposed spatial and temporal feature injection scheme enables tuning-free adaptation of I2V generation models for video editing. Our experimental results demonstrate that each component in our design is crucial to the accurate editing of source videos. We showcase more qualitative results for the effectiveness of our model components in Section 5.

### v3

@@ Abstract
In the dynamic field of digital content creation using generative models, state-of-the-art video editing models still do not offer the level of quality and control that users desire. Previous works on video editing either extended from image-based generative models in a zero-shot manner or necessitated extensive fine-tuning, which can hinder the production of fluid video edits. Furthermore, these methods frequently rely on textual input as the editing guidance, leading to ambiguities and limiting the types of edits they can perform. Recognizing these challenges, we introduce AnyV2V, a novel tuning-free paradigm designed to simplify video editing into two primary steps: (1) employing an off-the-shelf image editing model to modify the first frame, (2) utilizing an existing image-to-video generation model to generate the edited video through temporal feature injection. AnyV2V can leverage any existing image editing tools to support an extensive array of video editing tasks, including prompt-based editing, reference-based style transfer, subject-driven editing, and identity manipulation, which were unattainable by previous methods. AnyV2V can also support any video length. Our evaluation indicates that AnyV2V significantly outperforms other baseline methods in automatic and human evaluations by significant margin, maintaining visual consistency with the source video while achieving high-quality edits across all the editing tasks.
Figure 1: AnyV2V is an evolving framework to handle all types of video-to-video editing tasks without any parameter tuning. AnyV2V disentangles video editing into two simpler problems: (1) Single image editing and (2) Image-to-video generation with video reference.
@@ 1 Introduction
The development of deep generative models [23] has led to significant advancements in content creation and manipulation, especially in digital images [43, 36, 4, 28, 9, 30]. However, video generation and editing have not reached the same level of advancement as images [51, 6, 7, 24]. In the context of video editing, training a large-scale video editing model presents considerable challenges due to the scarcity of paired data and the substantial computational resources required.
To overcome these challenges, researchers proposed various approaches [16, 12, 52, 53, 32, 31, 19, 12, 60, 39, 5, 56, 27, 20], which can be categorized into two types: (1) zero-shot adaptation from pre-trained text-to-image (T2I) models or (2) fine-tuned motion module from a pre-trained T2I or text-to-video (T2V) models. The zero-shot methods [16, 12, 27, 60] usually suffer from flickering issues due to a lack of temporal understanding. On the other hand, fine-tuning methods [52, 8, 53, 19, 20] require more time and computational overhead to edit videos. Moreover, all the methods can only adhere to certain types of edits. For example, a user might want to perform edits on visuals that are out-of-distribution from the learned text encoder (e.g. change a person to a character from their artwork). Thus, a highly customizable solution is more desired in video editing applications. It would be ideal if there were methods that allowed for a seamless combination of human artistic input and the assistance provided by AI, allowing for a synergy between human effort and AI, maintaining creators’ originality and creativity while producing high-quality results.
With the above motivations, we aim to develop a video editing framework that requires no fine-tuning and caters to user demand. In this work, we present AnyV2V, designed to enhance the controllability of zero-shot video editing by decomposing the task into two pivotal stages:
@@
    Apply image edit on the first frame with any off-the-shelf image editing model or human effort.
@@
    Leverage the innate knowledge of the image-to-video (I2V) model to generate the edited video with the edited first frame, source video latent, and the intermediate temporal features.
@@
Our objective is to propagate the edited first frame across the entire video while ensuring alignment with the source video. To achieve this, we employ I2V models [59, 10, 42] for DDIM inversion to enable first-frame conditioning. With the inverted latents as initial noise and the modified first frame as the conditional signal, the I2V model can generate videos that are not only faithful to the edited first frame but also follow the appearance and motion of the source video. To further enforce the consistency of the appearance and motion with the source video, we perform feature injection in the I2V model. The two-stage editing process effectively offloads the editing operation to existing image editing tools. This design (detail in Section 3) helps AnyV2V excel in:
  • Compatibility: It provides a highly customized interface for a user to perform video edits on any modality. It can seamlessly integrate any image editing methods to perform diverse editing.
@@
  • Simplicity: It does not require any fine-tuning nor additional video features like previous works [19, 53, 38] to achieve high appearance and temporal consistency for video editing tasks.
From our findings, without any fine-tuning, AnyV2V can perform video editing tasks beyond the scope of current publicly available methods, such as reference-based style transfer, subject-driven editing, and identity manipulation. AnyV2V also can perform prompt-based editing and achieve superior results on common video editing evaluation metrics compared to the baseline models [16, 12, 53]. We show both quantitatively and qualitatively that our method outperforms existing SOTA baselines in Section 4.2 and Appendix B. AnyV2V is favoured in 69.7% of samples for prompt alignment and 46.2% overall preference in human evaluation, while the best baseline only achieves 31.7% and 20.7%, respectively. Our method also reaches the highest CLIP-Text score of 0.2932 in text alignment and a competitive CLIP-Image score of 0.9652 in temporal consistency.
All these achievements are thanks to the AnyV2V’s design to harness the power of off-the-shelf image editing models from advanced image editing research. Through a comprehensive study and evaluation of the effectiveness of our design, our key observation is that the inverted noise latent and feature injection serve as critical components to guide the video motion, and the I2V model itself has good capabilities in generating motions. We also found that by inverting long videos exceeding the I2V models’ training frames, the inverted latents enable the I2V model to produce longer videos, making long video editing possible. To summarize, The main contributions of our work are three-fold:
  • We proposed AnyV2V as a first fundamentally different solution for video editing, treating video editing as a simpler image editing problem.
  • We showed that AnyV2V can support long video editing by inverting videos that extend beyond the training frame lengths of I2V models.
  • Our extensive experiments showcased the superior performance of AnyV2V when compared to the existing SOTA methods, highlighting the potential of leveraging I2V models for video editing.
2 Related Works
Video generation has attracted considerable attention within the field [6, 7, 37, 51, 25, 57, 21, 50, 55, 10, 3, 42, 59]. However, video manipulation also represents a significant and popular area of interest. Initial attempts, such as Tune-A-Video [53] and VideoP2P [33] involved fine-tuning a text-to-image model to achieve video editing by learning the continuous motion. The concurrent works at that time such as Pix2Video [5] and Fate-Zero [39] go for zero-shot approach, which leverages the inverted latent from a text-to-image model to retain both structural and motion information. The progressively propagates to other frames edits. Subsequent developments have enhanced the results but generally follow the two paradigms  [16, 52, 12, 56, 5, 38, 20, 19, 14, 8, 27, 60, 11]. Control-A-Video [8] and ControlVideo [60] leveraged ControlNet [58] for extra spatial guidance. TokenFlow [16] leveraged the nearest neighbor field and inverted latent to achieve temporally consistent edit. Fairy [52] followed both paradigms which they fine-tuned a text-to-image model and also cached the attention maps to propagate the frame edits. VideoSwap [19] requires additional parameter tuning and video feature extraction (e.g. tracking, point correspondence, etc) to ensure appearance and temporal consistency. CoDeF [38] allows the first image edit to propagate the other frames with one-shot tuning. UniEdit [2] leverages the inverted latent and feature maps injection to achieve a wide range of video editing with a pre-trained text-to-video model [51].
Table 1: Comparison with different video editing methods and the type of editing tasks.
Method
Prompt-based
Editing
Reference-based
Style Transfer
Subject-Driven
Editing
Identity
Manipulation
Tuning-Free? Backbone
Tune-A-Video [53] Stable Diffusion [43]
Pix2Video [5] SD-Depth  [43]
Gen-1 [14] Stable Diffusion [43]
TokenFlow [16] Stable Diffusion [43]
FLATTEN [12] Stable Diffusion [43]
Fairy [52] Stable Diffusion [43]
ControlVideo [60] ControlNet [58]
CoDeF [38] ControlNet [58]
VideoSwap [19] AnimateDiff [20]
UniEdit [2] Any T2V [51]
AnyV2V (Ours) Any I2V [42, 59, 10]
However, none of the methods can offer precise control to users, as the edits may not align with the user’s exact intentions or desired level of detail, often due to the ambiguity of natural language and the constraints of the model’s capabilities. For example, VideoP2P [32] is restricted to only word-swapping prompts due to the reliance on cross-attention. There is a clear need for a more precise and comprehensive solution for video editing tasks. Our work AnyV2V is the first work to empower a diverse array of video editing tasks. We compare AnyV2V with the existing methods in Table 1. As can be seen, our method excels in its applicability and compatibility.
@@ 3 AnyV2V
Figure 2: AnyV2V takes a source video $V^{S}$ as input. In the first stage, we apply a block-box image editing method on the first frame $I_{1}$ according to the editing task. In the second stage, the source video is inverted to initial noise $z_{T}^{S}$, which is then denoised using DDIM sampling. During the sampling process, we extract spatial convolution, spatial attention, and temporal attention features from the I2V models’ decoder layers. To generate the edited video, we perform a DDIM sampling by fixing $z_{T}^{*}$ as $z_{T}^{T}$ and use the edited first frame as the conditional signal. During sampling, we inject the features and attention into corresponding layers of the model.
Our method presents a two-stage approach to video editing. Given a source video $V^{S}=\{I_{1},I_{2},I_{3},...,I_{n}\}$, where $I_{i}$ is the frame at time $i$ and $n$ denotes the video length, we first extract the initial frame $I_{1}$ and pass it into an image editing model $\phi_{\text{img}}$ to obtain an edited first frame $I^{*}_{1}=\phi_{\text{img}}(I_{1},C)$. $C$ is the auxiliary conditions for image editing models such as text prompt, mask, style, etc. In the second stage, we feed the edited first frame $I^{*}_{1}$ and a target prompt $\mathbf{s}^{*}$ into an I2V generation model $\epsilon_{\theta}$ and employ the inverted latent from the source video $V^{S}$ to guide the generation process such that the edited video $V^{*}$ follows the motion of the source video $V^{S}$, the semantic information represented in the edited first frame $I^{*}_{1}$, and the target prompt $\mathbf{s^{*}}$. An overall illustration of our video editing pipeline is shown in Figure 2. We provide a discussion of the preliminaries of our method in Appendix C. In this section, we explain each core component of our method.
@@ 3.1 Flexible First Frame Editing
In visual manipulation, controllability is a key element in performing precise editing. AnyV2V enables more controllable video editing by utilizing image editing models to modify the video’s first frame. This strategic approach enables highly accurate modifications in the video and is compatible with a broad spectrum of image editing models, including other deep learning models that can perform image style transfer [15, 17, 34, 48], mask-based image editing [36, 1], image inpainting [45, 29], identity-preserving image editing [49], and subject-driven image editing [9, 30, 18].
@@ 3.3 Appearance Guidance via Spatial Feature Injection
Our empirical observation (Section 4.3) suggests that I2V generation models already have some editing capabilities by only using the edited first frame and DDIM inverted noise as the model input. However, we find that this simple approach is often unable to correctly preserve the background in the edited first frame and the motion in the source video, as the conditional signal from the source video encoded in the inverted noise is limited.
To enforce consistency with the source video, we perform feature injection in both convolution layers and spatial attention layers in the denoising U-Net. During the video sampling process, we simultaneously denoise the source video using the previously collected DDIM inverted latents $\mathbf{z}^{S}_{t}$ at each time step $t$ such that $\mathbf{z}^{S}_{t-1}=\epsilon_{\theta}(\mathbf{z}^{S}_{t},I_{1},\varnothing,t)$. We preserve two types of features during source video denoising: convolution features $f^{l_{1}}$ before skip connection from the $l_{1}^{\text{th}}$ residual block in the U-Net decoder, and the spatial self-attention scores $\{A_{s}^{l_{2}}\}$ from $l_{2}=\{l_{low},l_{low+1},...,l_{high}\}$ layers. We collect the queries $\{Q_{s}^{l_{2}}\}$ and keys $\{K_{s}^{l_{2}}\}$ instead of directly collecting $A_{s}^{l_{2}}$ as the attention score matrices are parameterized by the query and key vectors. We then replace the corresponding features during denoising the edited video in both the normal denoising branch and the negative prompt branch for classifier-free guidance [22]. We use two thresholds $\tau_{conv}$ and $\tau_{sa}$ to control the convolution and spatial attention injection to only happen in the first $\tau_{conv}$ and $\tau_{sa}$ steps during video sampling.
@@ 3.4 Motion Guidance through Temporal Feature Injection
To better reconstruct the source video motion in the edited video, we propose to also inject the temporal attention features in the video generation process. Similar to spatial attention injection, we collect the source video temporal self-attention queries $Q^{l_{3}}_{t}$ and keys $K^{l_{3}}_{t}$ from some U-Net decoder layers represented by $l_{3}$ and inject them into the edited video denoising branches. We also only apply temporal attention injection in the first $\tau_{ta}$ steps during sampling.
3.5 Putting it Together
Overall, combining the spatial and temporal feature injection mechanisms, we replace the editing branch features $\{f^{*l_{1}},Q_{s}^{*l_{2}},K_{s}^{*l_{2}},Q_{t}^{*l_{3}},K_{t}^{*l_{3}}\}$ with the features from the source denoising branch:
$\mathbf{z}^{*}_{t-1}=\epsilon_{\theta}(\mathbf{z}^{*}_{t},I^{*},\mathbf{s}^{*},t\ ;\{f^{l_{1}},Q_{s}^{l_{2}},K_{s}^{l_{2}},Q_{t}^{l_{3}},K_{t}^{l_{3}}\}),$ (2)
where $\epsilon_{\theta}(\cdot\ ;\{f^{l_{1}},Q_{s}^{l_{2}},K_{s}^{l_{2}},Q_{t}^{l_{3}},K_{t}^{l_{3}}\})$ denotes the feature replacement operation across different layers $l_{1},l_{2},l_{3}$. Our proposed spatial and temporal feature injection scheme enables tuning-free adaptation of I2V generation models for video editing. Our experimental results demonstrate that each component in our design is crucial to the accurate editing of source videos. We showcase more qualitative results for the effectiveness of our model components in Section 4.